# Collaborative Learning in Agentic Systems: A Collective AI is Greater Than the Sum of Its Parts

### v1

@@ Abstract
Agentic AI has gained significant interest as a research paradigm focused on autonomy, self-directed learning, and long-term reliability of decision making. Real-world agentic systems operate in decentralized settings on a large set of tasks or data distributions with constraints such as limited bandwidth, asynchronous execution, and the absence of a centralized model or even common objectives. We posit that exploiting previously learned skills, task similarities, and communication capabilities in a collective of agentic AI are challenging but essential elements to enabling scalability, open-endedness, and beneficial collaborative learning dynamics. In this paper, we introduce Modular Sharing and Composition in Collective Learning (MOSAIC), an agentic algorithm that allows multiple agents to independently solve different tasks while also identifying, sharing, and reusing useful machine-learned knowledge, without coordination, synchronization, or centralized control. MOSAIC combines three mechanisms: (1) modular policy composition via neural network masks, (2) cosine similarity estimation using Wasserstein embeddings for knowledge selection, and (3) asynchronous communication and policy integration. Results on a set of RL benchmarks show that MOSAIC has a greater sample efficiency than isolated learners, i.e., it learns significantly faster, and in some cases, finds solutions to tasks that cannot be solved by isolated learners. The collaborative learning and sharing dynamics are also observed to result in the emergence of ideal curricula of tasks, from easy to hard. These findings support the case for collaborative learning in agentic systems to achieve better and continuously evolving performance both at the individual and collective levels.
@@ 1 Introduction
The ability to handle multiple dynamic tasks concurrently and over long-term time windows De Lange et al. (2022); Parisi et al. (2019) is a coveted feature of autonomous, adaptive agents that operate in open-ended environments involving physical-, vision-, and language-based tasks. Dynamic environments demand resilient agents that continuously learn useful behaviors from local data, generalize across diverse tasks, and effectively share useful knowledge (Tarale et al., 2025). Such requirements are central to the emerging vision of agentic AI systems, which emphasize self-directed learning, adaptation, asynchronous operation, and scalable autonomy (Colas et al., 2022; Forestier et al., 2022). Recent work on collaborative and collective AI further highlights the importance of decentralized learning, peer-to-peer sharing, and modular reuse for building scalable societies of AI agents (Soltoggio et al., 2024).
In such scenarios, reinforcement learning (RL) signals that can be sourced from the environment allow for policy optimization without external supervision, making RL an effective tool used both in robotics and language models (Sutton and Barto, 2018; Jaderberg et al., 2017; Wang et al., 2025; Sun, 2023). Crucially, knowledge reuse has been a central assumption in multi-task (MTRL) and meta reinforcement learning (Meta-RL) (Hospedales et al., 2021), in which learning of many tasks is facilitated by exploiting similarities across tasks. The successful deployment of pre-trained models in LLMs is another example of effective reuse of previously acquired knowledge across tasks Han et al. (2021); Qiu et al. (2020). Reliability of RL algorithms when dealing with evolving streams of data is improved with lifelong RL approaches (Ben-Iwhiwhu et al., 2023).
Most existing MTRL, Meta-RL, and lifelong RL methods assume a centralized setup with shared policy networks and known task labels (Wang et al., 2023; Sun et al., 2022; Hendawy et al., 2024; Amani et al., 2023; Zeng et al., 2021; D’Eramo et al., 2020). These assumptions limit their applicability in fully decentralized or real-world agentic scenarios, where agents typically operate independently, and communication, when possible, is asynchronous and not coordinated. To address these realistic conditions, recent work in lifelong learning and multi-agent systems explored machine-learned knowledge sharing among many agents (Nath et al., 2023; Rostami et al., 2017; Gao et al., 2021; Ge et al., 2023; Gerstgrasser et al., 2023). Similar objectives are pursued by recent approaches that share and compose LoRA modules in LLMs (Huang et al., 2024; Roy et al., 2025), demonstrating that this challenge is crucial to scale current AI approaches. A full coverage of the related works in the literature can be found in appendix A.
While the idea of sharing and reusing knowledge in agentic machine learning systems is intuitive and appealing, a key algorithmic problem has not yet been well addressed: in a collective of AI agents solving different tasks, what specific knowledge should be shared, with whom, and when? DARPA/MTO (2021); Soltoggio et al. (2024) Moreover, does a decentralized collective of agents have an advantage over isolated agents learning the same tasks?
To address these unanswered questions, this study proposes a new algorithm, Modular Sharing and Composition in Collective Learning (MOSAIC), that describes what policies to exchange among peers, specifically among which agents, when to do so, and how such policies can be integrated. Extending and integrating recent advances in lifelong machine learning, the approach employs modular knowledge representations in the form of network masks and proposes a knowledge exchange mechanism based on cosine similarity computed with Wasserstein task embeddings. Performance measures and similarities are used by individual agents to determine specifically what knowledge to retrieve, from whom, and when.
The experimental evidence and the ablation studies show that precise selection of specific knowledge is essential to exploit the advantages of sharing. This evidence indicates that not all knowledge is beneficial to all agents, and that modular composition and cosine similarity metrics are essential ingredients in a collective of collaborating agents. MOSAIC agents learn faster than isolated agents by selectively leveraging the available knowledge from peers working on similar tasks. In some cases, by combining knowledge retrieved from other agents working on similar tasks, MOSAIC agents are able to solve tasks that are too difficult for isolated learners. Further analysis reveals that the collective implicitly discovers an ideal curriculum in which agents working on simpler tasks subsequently help agents working on harder tasks. The MOSAIC learning dynamics indicate that a collective of collaborative MOSAIC agents can achieve superior performance than that of an equivalent number of isolated learners, making a strong case for targeted knowledge sharing in agentic systems.
1.1 Contributions
This work introduces MOSAIC, an agentic system, i.e., an autonomous decision-making and self-contained individual agent, that implements autonomous and self-directed knowledge exchange with other MOSAIC peers. MOSAIC is a novel integration of the following algorithmic steps: (i) modularized knowledge learning and representation; (ii) online Wasserstein task embeddings computed from reinforcement learning datasets; (iii) knowledge querying and retrieval with cosine similarity (between embeddings) and performance-based selection heuristics.
The performance evaluations reveal significant learning advantages for a collective of collaborating agents with respect to isolated learners, which, in some cases, results in successful policies that cannot be discovered in isolation.
Lastly, the analysis of learning dynamics revealed that agents choose what to learn in ways that lead to the emergence of implicit curricula, task clusters, and task hierarchies. This property enables an open-ended training behavior in which agents autonomously determine their learning trajectories by reusing specific subsets from the available collective knowledge.
@@ 2 Methodology
The following section describes the components of the MOSAIC algorithm that includes knowledge representations (Section 2.1), Wasserstein task embedding (Section 2.2), knowledge selection and sharing (Section 2.3), and knowledge composition (Section 2.4). Finally, Section 2.5 illustrates the integration of all components with pseudo-code. A visual summary of the main algorithmic steps is provided in Figure 1.
@@ 2.1 Policy Representations
The ability to encapsulate specific knowledge in isolated module representations (Alet et al., 2018) is a key component of MOSAIC. To this aim, a variety of parameter isolation mechanisms from the field of lifelong learning can be used von Oswald et al. (2020); Hu et al. (2022); Rusu et al. (2016); Mallya and Lazebnik (2018). Among those, promising results in lifelong RL were shown by network masks (Ben-Iwhiwhu et al., 2023; Wortsman et al., 2020) that are therefore used in this study. Masks are naturally suited for composition through linear combinations, which is an essential mechanism in MOSAIC. The use of masking makes MOSAIC agnostic to the specific reinforcement learning algorithm used, requiring only that policies are parameterized by a neural network. This allows it to be layered on top of any standard RL method, such as PPO (Schulman et al., 2017), DQN (Mnih et al., 2013), or A3C (Mnih et al., 2016).
Figure 1: High-level illustration of the main MOSAIC algorithmic steps. (A) A Wasserstein task embedding $v_{\tau}$ is computed from a SAR batch (Section 2.2) by Agent A0 (representing any agent in the collective). (B) Periodically, the agent A0 broadcasts a task embedding query (TEQ) to all known peers. (C) Peers send back a query response (QR) that contains their task’s $v_{\tau},r$, and the corresponding mask ID; (D) Agent A0 selects relevant embeddings using cosine similarity on the Wasserstein embeddings. (E) A0 further selects relevant embeddings using the performance criterion (Section 2.3). (F) A0 sends mark requests (MR). (G) The contacted agents respond by sending the requested mask through a mask transfer (MTR). (H) Incoming masks from A2 and A3 are incorporated into A0 (Section 2.4). The training of the agent’s policy occurs in parallel (not represented in the figure).
Preliminaries.
Each task $\tau\in\mathcal{T}$ is modeled as an MDP defined by the tuple $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma\right\rangle$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}(s^{\prime}|s,a)$ is the transition probability function, $\mathcal{R}(s,a)\in\mathbb{R}$ is the reward received after taking action $a$ in the state $s$, and $\gamma\in[0,1]$ is the discount factor. An agent interacts with its environment and aims to learn a policy $\pi_{\tau}:\mathcal{S}\to\mathcal{A}$ that maximizes the expected discounted return,
$J_{\tau}(\pi_{\tau})=\mathbb{E}_{\pi_{\tau}}\left[\sum_{t=0}^{\infty}\gamma^{t}_{\tau}\mathcal{R}_{\tau}(s_{t},a_{t})\right].$ (1)
Mask Representations.
The mask representation used in this study relies on a frozen backbone network $\Phi$ that must be the same for all agents. Each policy $\pi_{\tau}$ is implicitly parameterized through a sparse mask $\phi_{\tau}$, such that $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau})}$, where $\odot$ denotes the Hadamard product, and $g(\cdot)$ is a binarization function applied during the forward pass. The mask $\phi_{\tau}$ consists of a set of real-valued score vectors, one per layer of the shared backbone network $\Phi$, which select a sparse, task-specific subnetwork from $\Phi$ without modifying its parameters. The scores are binarized during the forward pass to produce a hard mask that is used for inference. The binarization is performed during inference, using the element-wise thresholding function,
@@
where $g(\phi_{\tau})$ denotes the full binary mask and $\ell$ indexes the layers of the backbone network $\Phi$. During the backward pass, gradients are computed with respect to the real-valued mask parameters $\phi_{\tau}$, and updates are performed using the straight-through estimator (STE) (Wortsman et al., 2020) to enable learning through the non-differentiable binarization step.
@@ 2.2 Wasserstein Task Embeddings for Online Reinforcement Learning
A key intuition in this study is that knowledge composition is effective if policies from other agents are acquired only when they are likely to be beneficial, and because they have been derived from related tasks. This hypothesis was verified with an ablation study. MOSAIC implements this idea by computing task embeddings from trajectory data and using them to identify semantically related peers via geometric proximity. This idea builds on work in latent task inference and representation learning (Zintgraf et al., 2021; Achille et al., 2019), adopting a non-parametric approach aligned with prototype-based and optimal transport methods (Yarats et al., 2021; Chu et al., 2024; Liu et al., 2022).
Unlike gradient-based adaptation techniques (Lan et al., 2019), which are unstable in RL and costly at inference time, MOSAIC computes task embeddings through training-free distributional comparisons, making them well-suited for decentralized, real-time systems requiring modular policy reuse without supervision. To support these requirements, MOSAIC extends the Wasserstein Task Embedding (WTE) framework of Liu et al. (2022) to the RL domain by defining an online embedding function over batches of state-action-reward (SAR) tuples drawn from an agent’s replay buffer. For task $\tau$, the agent maintains an empirical task distribution $\mu_{\tau}\in\mathcal{P}(\mathbb{R}^{d})$, defined as,
$\mu_{\tau}=\frac{1}{N}\sum_{t=1}^{N}\delta_{x_{t}},\quad\text{where}\quad x_{t}=(s_{t},a_{t},r_{t})\quad\text{and}\quad x_{t}\sim\mathcal{D}_{\tau},$ (3)
where $\mathcal{D}_{\tau}$ denotes the replay buffer for task $\tau$, and each $x_{t}\in\mathbb{R}^{d}$ is a state-action-reward (SAR) tuple. The tuple dimensionality is given by $d=d_{s}+d_{a}+d_{r}$, where $d_{s}$, $d_{a}$, and $d_{r}$ are the dimensions of the state, action, and scalar reward, respectively. In our experiments, we use $N=128$ samples to compute the empirical task distribution. $\delta_{x_{t}}$ denotes the Dirac measure centered at the sample $x_{t}$. A fixed synthetic reference distribution $\mu_{0}$ is defined as,
@@
$\gamma^{*}=\arg\min_{\gamma\in\mathbb{R}^{N\times M}}\sum_{t=1}^{N}\sum_{m=1}^{M}\gamma_{tm}\|x_{t}-x^{0}_{m}\|^{2},\quad\text{s.t. }\gamma_{tm}\geq 0,\gamma\mathbf{1}_{M}=\frac{1}{N}\mathbf{1}_{N},\gamma^{\top}\mathbf{1}_{N}=\frac{1}{M}\mathbf{1}_{M},$ (5)
@@
This mapping places each task $\tau$ in a shared latent space, where metrics can be used to quantify task relationships. A 2-point moving average is used to smooth behavioral fluctuations and stabilize evolving policies in MOSAIC. This embedding strategy enables agents to reason about task-relatedness in constant time, without additional training or forward passes. Importantly, it fits the online and asynchronous nature of agentic RL systems, while remaining robust to evolving policies and lightweight enough for continuous use in edge-deployed systems.

### v2

# Policy Search, Retrieval, and Composition via Task Similarity in Collaborative Agentic Systems

@@ Abstract
Agentic AI aims to create systems that set their own goals, adapt proactively to change, and refine behavior through continuous experience. Recent advances suggest that, when facing multiple and unforeseen tasks, agents could benefit from sharing machine-learned knowledge and reuse policies that have already been fully or partially learned by other agents. However, how to query, select, and retrieve policies from a pool of agents, and how to integrate such policies remains a largely unexplored area. This study explores how an agent decides what knowledge to select, from whom, and when and how to integrate it in its own policy in order to accelerate its own learning. The proposed algorithm, Modular Sharing and Composition in Collective Learning (MOSAIC), improves learning in agentic collectives by combining (1) knowledge selection using performance signals and cosine similarity on Wasserstein task embeddings, (2) modular and transferable neural representations via masks, and (3) policy integration, composition and fine-tuning. MOSAIC outperforms isolated learners and global sharing approaches in both learning speed and overall performance, and in some cases solves tasks that isolated agents cannot. The results also demonstrate that selective, goal-driven reuse leads to less susceptibility to task interference. We also observe the emergence of self-organization, where agents solving simpler tasks accelerate the learning of harder ones through shared knowledge.
@@ 1 Introduction
The ability to solve complex, open-ended problems is desirable objective as agentic AI systems become increasingly central in AI architectures. In real-world open-ended settings, these systems face challenges in generalizing and adapting to an ever evolving stream of data and tasks (De Lange et al. 2022; Parisi et al. 2019; Lesort et al. 2020). While new lifelong learning algorithms (Kudithipudi et al. 2022) are being introduced to provide continual evolution and adaptation, when agents learn in isolation, they can only benefit from their limited experiences, unlike humans who benefit from collaboration and sharing experiences and skills. Recent studies suggest that a shift toward more collaborative types of learning, where agents selectively share and reuse modular knowledge from other agents, is key to reduce redundant learning, accelerate adaptation, and improve robustness (Tarale, Rietman, and Siegelmann 2025; Soltoggio et al. 2024; DARPA/MTO 2021).
In the context of self-adaptation, reinforcement signals provide an effective tool to achieve continuous improvement in a wide range of domains (Sutton and Barto 2018; Jaderberg et al. 2017; Wang et al. 2025; Sun 2023). In addition, task similarities have been exploited to benefit learning in areas such as transfer learning and multi-task learning (Hendawy, Peters, and D’Eramo 2024; Wang et al. 2023; Sun et al. 2022; D’Eramo et al. 2020; Zeng et al. 2021; Espeholt et al. 2018). Lifelong learning approaches also leverage task similarity when previously learned tasks benefit the learning of new tasks, with an advantage often named forward transfer (Ben-Iwhiwhu et al. 2023; Kirkpatrick et al. 2017; Wolczyk et al. 2022; Kessler et al. 2022; Cannon and Şimşek 2024). However, many approaches to share knowledge, e.g., distributed RL (Sartoretti et al. 2019) or Federated Learning (Wen et al. 2023), often assume some form of centralization and uniformity of tasks, which is ill-suited to the requirements of agentic AI where each agents acts independently, asynchronously, and likely on a large variety of different tasks. Recent studies have begun to investigate sharing and reuse of policies in less constrained, evolving multi-agent contexts (Nath et al. 2023; Tarale, Rietman, and Siegelmann 2025; Rostami et al. 2017; Amani et al. 2023; Gao et al. 2021; Gerstgrasser, Danino, and Keren 2023). These developments reflect a growing recognition that scalable AI systems increasingly depend on selective, self-directed collaboration among autonomous learners.
Knowledge and skill sharing among a set of independent and autonomous learners, while conceptually appealing, presents significant challenges. In particular, open research questions reflect the problems of how to identify what knowledge to acquire, from whom to acquire it, how such knowledge is best represented and and how to integrate it to manifest robust lifelong learning properties (DARPA/MTO 2021; Soltoggio et al. 2024). Furthermore, it is still unclear to what extent collaborative learning offers an advantage over isolated or centralized learning.
To address these questions, this study introduces Modular Sharing and Composition in Collective Learning (MOSAIC) [^1] An implementation of MOSAIC is available at https://github.com/DMIU-ShELL/MOSAIC., a novel algorithm that proposes a solution to select appropriate knowledge across a collective of agents with the aim of implementing policy composition and reuse. The problem analysis and following ablation studies indicate that essential components of such a system are (1) knowledge selection via task similarity, (2) modular and transferable neural representations, and (3) policy integration, composition and fine-tuning.
Specific algorithmic choices in MOSAIC to implement such components are Wasserstein embeddings for task similarity, transferable binary network masks for sharing, and learnable linear combinations of those to achieve integration and fine tuning. Alternative but equivalent methods could also be used to adapt MOSAIC to a variety of domains, e.g., LoRA-based modular composition for foundation models and LLMs (Hu et al. 2022; Huang et al. 2024; Zhong et al. 2024; Asadi et al. 2024; Prabhakar et al. 2025), and embedding methods or descriptors to express task alignment (Duan et al. 2017; Achille et al. 2019; Dou, Luo, and Yang 2022; Grover et al. 2018; Rakelly et al. 2019).
Simulations indicate that selectively acquiring relevant knowledge is essential to effectively reuse policies. The composition of policies is best achieved with a similarity-driven and reward-aware weighting. The results show how communicating MOSAIC agents learn faster than isolated learners by reusing knowledge from peers on similar tasks. In some cases, communicating agents are able to solve tasks that they could not learn alone. In addition, the analysis reveals an implicit self-organization of learning into task curricula, where agents completing simpler tasks accelerate the progress of those tackling harder ones.
In summary, MOSAIC supports the hypothesis that task-relevant policy transfer and reuse, guided by agent-centered optimization objectives, is feasible and effective with advantages over learning in isolation or sharing global parameters, which could lead to task interference. To the best of our knowledge, this is the first study that combines modular and transferable task-specific knowledge with similarity-based selection and integration. Such a combination of components are sufficient to significantly improve sample efficiency and learning through collective knowledge sharing.
2 Related Work
Recent work in modular knowledge reuse for multi-agent RL explores various approaches to enable efficient knowledge transfer. Federated learning methods (Yoon et al. 2021; Seo, Han, and Yoo 2025) communicate task-specific masks or binary representations. These approaches lack principled task similarity metrics for selective reuse. MOSAIC addresses this gap through Wasserstein-based similarity measures that guide knowledge selection based on performance signals.
Several frameworks enable parameter-efficient knowledge sharing through masking (Paul et al. 2025; Gao et al. 2021; Ge et al. 2023; Nath et al. 2023; Rostami et al. 2017; Gerstgrasser, Danino, and Keren 2023), prototype aggregation (Tan et al. 2022), or decentralized training (Douillard et al. 2024; Jaghouar, Ong, and Hagemann 2024). Communication-efficient methods reduce information exchange through knowledge preservation and selective sharing (Diskin et al. 2021; McMahan et al. 2017; Meng et al. 2025). These approaches rely on predefined sharing protocols or require synchronous updates. MOSAIC’s binary mask representation enables asynchronous, selective reuse and maintains modularity while minimizing communication overhead.
Policy composition methods like PaCo (Sun et al. 2022) interpolate within shared subspaces. These methods lack mechanisms for similarity-driven integration. Peer-to-peer frameworks (Sengupta et al. 2024; Alvar and Bajić 2019; Kaimakamidis, Mademlis, and Pitas 2024; Tarale, Rietman, and Siegelmann 2025; Yu, Vincent, and Schwager 2022; Nadiger, Kumar, and Abdelhak 2019; Lozano-Cuadra et al. 2025) and model-based approaches (Jiang, Narayanan, and Li 2021; Du et al. 2025; Shi et al. 2022) enable decentralized coordination. These frameworks do not combine performance-based selection with modular representations. MOSAIC integrates Wasserstein-based task similarity with binary mask modularity and similarity-driven linear policy combinations. MOSAIC demonstrates improved sample efficiency and self-organization in collaborative multi-agent systems. Further discussion of related works is provided in Appendix A.
@@ 3 Methodology
The following section describes the components of the MOSAIC algorithm. A high level illustration is provided in Figure 1. Pseudocode can be found in the Appendix E.
@@ 3.1 Policy Representations
Isolating task-specific knowledge in compact representations (Alet, Lozano-Perez, and Kaelbling 2018) is a key to make policies easy to transfer across agents. Several parameter isolation mechanisms from lifelong learning can achieve this goal (von Oswald et al. 2020; Hu et al. 2022; Rusu et al. 2016; Mallya and Lazebnik 2018).
This study adopts neural network masks (Ben-Iwhiwhu et al. 2023; Wortsman et al. 2020) that have shown strong results in lifelong reinforcement learning and supervised learning settings, and supports composition through linear combinations. This design enables integration with standard methods such as PPO (Schulman et al. 2017), DQN (Mnih et al. 2013), or A3C (Mnih et al. 2016).
Figure 1: High-level illustration of the main MOSAIC algorithmic steps. (A) A Wasserstein task embedding $v_{\tau}$ is computed from a SAR batch by Agent A0 (representing any agent in the collective). (B) Periodically, the agent A0 broadcasts a task embedding query (TEQ) to all known peers. (C) Peers send back a query response (QR) that contains their task’s $v_{\tau},r$, and the corresponding mask ID; (D) Agent A0 selects relevant embeddings using cosine similarity on the Wasserstein embeddings. (E) A0 further selects relevant embeddings using Criterion 1 and Criterion 2. (F) A0 sends mark requests (MR). (G) The contacted agents respond by sending the requested mask through a mask transfer (MTR). (H) Incoming masks from A2 and A3 are incorporated into A0. The training of the agent’s policy occurs in parallel (not represented in the figure).
Mask representations rely on a frozen backbone network $\Phi$ that is homogeneous across all agents. Each policy $\pi_{\tau}$ is implicitly parameterized through a sparse mask $\phi_{\tau}$, such that $\pi_{\tau}=\pi_{\Phi\odot g(\phi_{\tau})}$, where $\odot$ denotes the Hadamard product, and $g(\cdot)$ is a binarization function applied during the forward pass. The mask $\phi_{\tau}$ consists of a set of real-valued score vectors, one per layer of the shared backbone network $\Phi$, which select a sparse, task-specific subnetwork from $\Phi$ without modifying its parameters. Scores are binarized during the forward pass to select a discrete subnetwork from the backbone. The binarization is given by the element-wise thresholding function,
@@
where $g(\phi_{\tau})$ denotes the full binary mask and $\ell$ indexes the layers of the backbone network $\Phi$. During the backward pass, gradients are computed with respect to the real-valued mask parameters $\phi_{\tau}$, and updates are performed using the straight-through estimator (STE) (Wortsman et al. 2020) to enable learning through the non-differentiable binarization step.
@@ 3.2 Wasserstein Task Embeddings for Online Reinforcement Learning
To identify suitable policies for a given tasks, MOSAIC extends the Wasserstein Task Embedding (WTE) framework of Liu et al. (2025) to RL by computing embeddings over batches of state-action-reward (SAR) data collected online (see Figure 1(A)). For task $\tau$, the agent maintains an empirical task distribution
$\mu_{\tau}=\frac{1}{N}\sum_{t=1}^{N}\delta_{x_{t}}\in\mathcal{P}(\mathbb{R}^{d}),$ (2)
where $x_{t}=(s_{t},a_{t},r_{t})$ and $x_{t}\sim\mathcal{D}_{\tau}$. Here, $\mathcal{D}_{\tau}$ denotes the replay buffer for task $\tau$, and each $x_{t}\in\mathbb{R}^{d}$ is a state-action-reward (SAR) tuple. The tuple dimensionality is $d=d_{s}+d_{a}+d_{r}$, where $d_{s}$, $d_{a}$, and $d_{r}$ are the dimensions of the state, action, and scalar reward, respectively. We use $N=128$ samples in our experiments. $\delta_{x_{t}}$ denotes the Dirac measure centered at the sample $x_{t}$. A fixed synthetic reference distribution $\mu_{0}$ is defined as,
@@
$\gamma^{*}$ $=\arg\min_{\gamma\in\mathbb{R}^{N\times M}}\sum_{t=1}^{N}\sum_{m=1}^{M}\gamma_{tm}\|x_{t}-x^{0}_{m}\|^{2}$ (4)
$\quad\text{s.t. }\gamma_{tm}\geq 0,\;\gamma\mathbf{1}_{M}=\tfrac{1}{N}\mathbf{1}_{N},\;\gamma^{\top}\mathbf{1}_{N}=\tfrac{1}{M}\mathbf{1}_{M},$
@@
This mapping places each task $\tau$ in a shared latent space, where metrics can be used to quantify task relationships.